# GeoEdit: Geometric Knowledge Editing for Large Language Models

### v1

@@ Abstract
Regular updates are essential for maintaining up-to-date knowledge in large language models (LLMs). Consequently, various model editing methods have been developed to update specific knowledge within LLMs. However, training-based approaches often struggle to effectively incorporate new knowledge while preserving unrelated general knowledge. To address this challenge, we propose a novel framework called Geometric Knowledge Editing (GeoEdit). GeoEdit utilizes the geometric relationships of parameter updates from fine-tuning to differentiate between neurons associated with new knowledge updates and those related to general knowledge perturbations. By employing a direction-aware knowledge identification method, we avoid updating neurons with directions approximately orthogonal to existing knowledge, thus preserving the model’s generalization ability. For the remaining neurons, we integrate both old and new knowledge for aligned directions and apply a “forget-then-learn” editing strategy for opposite directions. Additionally, we introduce an importance-guided task vector fusion technique that filters out redundant information and provides adaptive neuron-level weighting, further enhancing model editing performance. Extensive experiments on two publicly available datasets demonstrate the superiority of GeoEdit over existing state-of-the-art methods.
@@ 1 Introduction
Large language models (LLMs) have demonstrated the ability to store vast amounts of knowledge during pre-training and retrieve it during inference Yao et al. (2023); Wang et al. (2024e). However, much of the knowledge in the real world is constantly evolving. For instance, the answer to the question “Who is the President of the United States?” was “Joe Biden” in 2024, but it is now “Donald Trump”. As a result, some knowledge that was once correct in LLMs can become obsolete or inaccurate Li and Chu (2024); Huang et al. (2024).
To address this issue, model editing methods have been developed to update the target new knowledge while preserving unrelated general knowledge within the model Hong and Lipani (2024); Ma et al. (2024); Wang et al. (2024f). Specifically, current model editing methods typically follow the “locate-and-edit” paradigm Wang et al. (2023, 2024a); Li et al. (2024). The core idea is first to locate influential weights in LLMs and then edit them by introducing a perturbation. Although effective, these approaches incur significant computational overhead to identify the important neurons and parameters Meng et al. (2022). Some methods also require sampling additional data (e.g., from Wikipedia) to mitigate the impact on the general knowledge within LLMs during editing Meng et al. (2023); Fang et al. (2024), introducing extra costs and potential biases.
@@
Additionally, to mitigate angular bias in high-dimensional space, GeoEdit projects the task vectors into a lower-dimensional semantic space using an auto-encoder. Angle calculations and editing operations are performed in this compressed space, after which the decoder reconstructs the final editing task vector $\tau_{edit}$ for parameter updates. To optimize vector fusion, we introduce an importance-guided task vector fusion technique, which applies fine-grained weights to the vectors and suppresses noise from redundant parameters, further enhancing the effectiveness of model editing.
@@ 2 Related Work
Knowledge editing has gained significant attention due to the increasing need to update the knowledge in LLMs Wang et al. (2024e); Bi et al. (2024).
Existing methods can be classified as follows: (1) Fine-tuning is an intuitive and straightforward way to update the model’s knowledge Feng et al. (2023); Gangadhar and Stratos (2024); Zheng et al. (2024). Recent parameter-efficient fine-tuning methods like Prefix-Tuning Li and Liang (2021) and LoRA Hu et al. (2021) have made fine-tuning more suitable for knowledge editing. Ni et al. (2024) proposes the “Forgetting before Learning” (F-Learning) paradigm, which employs parametric arithmetic to facilitate the forgetting of old knowledge and the learning of new knowledge. However, this often leads to unintended changes in unrelated knowledge, negatively affecting generalization Wang et al. (2024b); Gu et al. (2024). (2) Meta-learning (“learning to learn”) methods use a hypernetwork to generate task-specific updates, allowing LLMs to quickly adapt to new data without retraining from scratch De Cao et al. (2021); Shi et al. (2024); Wang et al. (2024d). (3) Locate-and-edit methods usually locate influential parameters and then edit them by introducing a perturbation Zhang et al. (2024); Jiang et al. (2024); Xu et al. (2024). Classic methods like ROME Meng et al. (2022) and MEMIT Meng et al. (2023) use causal reasoning to identify key neuron activations and adjust specific weights. More recent approaches, such as AlphaEdit Fang et al. (2024), improve the method by projecting perturbations onto the null space of preserved knowledge.
@@
However, fine-tuning and locate-and-edit methods struggle to balance new knowledge updates with preserving unrelated knowledge Gupta et al. (2024); Feng et al. (2024a). For instance, locate-and-edit methods typically require large additional datasets to capture general knowledge and avoid disruption during editing Wang et al. (2024c); Hsueh et al. (2024). However, these approaches introduce extra costs and biases.
@@
Thus, our paper focuses on improving fine-tuning methods. Motivated by F-Learning Ni et al. (2024), we propose distinguishing between general knowledge and updated knowledge based on the angular divergence between the updated directions of old and new knowledge. This allows us to avoid updating general knowledge, ensuring model generalization while applying tailored strategies to improve the effectiveness of knowledge updates.
3 Preliminary
@@
Figure 2: Overview of GeoEdit. Step (a): Neuron-level task vectors $\tau_{old}$ and $\tau_{new}$ are extracted for both the old and new knowledge datasets using parametric arithmetic. Step (b): An auto-encoder is trained to project a low-dimensional representation of the task vectors, eliminating the angular bias issue in high-dimensional space. Step (c): The latent task vectors, $h_{new}$ and $h_{old}$, are used to classify neurons based on the angle between their corresponding task vectors, followed by the application of different editing strategies to obtain $h_{edit}$. Finally, $h_{edit}$ is passed through the decoder to produce $\tau_{edit}$, which is added to the initial model to generate the edited model $f_{\theta_{e}}$.
Task Definition
@@ 4.1 Extracting Neuron-level Task Vectors
While prior research typically captures task vectors at the model level Ilharco et al. (2022), we propose extracting them at the neuron level for finer control. Let $\theta=\{\theta^{1},\theta^{2},\dots,\theta^{N}\}$ represent the $N$ neurons in the LLM, where the $i$-th neuron is represented by $\theta^{i}\in\mathbb{R}^{d_{n}}$ with $d_{n}$ dimensional parameters. The neuron-level task vectors are given by $\tau_{new}=\{\tau_{new}^{1},\tau_{new}^{2},\dots,\tau_{new}^{N}\}$, where $\tau_{new}^{i}$ corresponds to the new knowledge task vector for the $i$-th neuron. This approach enables more granular analysis of parameter changes and selective editing of knowledge-specific neurons, enhancing model editing precision.
@@
4.2 Dimensionality Reduction Using AE
After obtaining task vectors $\tau_{old}$ and $\tau_{new}$, we use an auto-encoder (AE) to map them into a low-dimensional latent space for knowledge editing. This step mitigates the angular bias inherent in high-dimensional task vectors. For example, in LLaMA-7B, each neuron-level task vector has 4096 dimensions, which naturally tend toward orthogonality, distorting the interpretation of their relationships. Dimensionality reduction also removes irrelevant information, improving the editing accuracy.
Thus, we train a semantic encoder and decoder, both implemented using multi-layer perceptrons (MLPs). The encoder maps high-dimensional task vectors to a lower-dimensional space for editing, while the decoder reconstructs the edited vectors, which are then added back to the initial model for the final results. Specifically, the semantic encoder, denoted as SemEnc($\cdot$), maps the high-dimensional task vectors $\tau_{old}$ and $\tau_{new}$ into the latent space as:
@@ 4.3 Geometric Knowledge Editing
After training the auto-encoder, we project $\tau_{old}$ and $\tau_{new}$ into the low-dimensional latent space for knowledge editing. GeoEdit then edit the latent task vectors to obtain $h_{edit}$, which is decoded to produce the edited task vector $\tau_{edit}$. This vector is subsequently added to the initial model, resulting in the final edited model $f_{\theta_{e}}$.
@@
For neuron $i$, we first use the encoder to reduce the dimensionality of $\tau_{old}^{i}$ and $\tau_{new}^{i}$ yielding the latent task vectors $h_{old}^{i}$ and $h_{new}^{i}$. We then compute the angular divergence $\phi$ between $h_{old}^{i}$ and $h_{new}^{i}$ as:
$\phi=\arccos\frac{h_{old}^{i}\cdot h_{new}^{i}}{\left|h_{old}^{i}\right|\cdot\left|h_{new}^{i}\right|}$ (10)
@@
$h_{edit}^{i}=\begin{cases}\alpha^{i}h_{old}^{i}+\beta^{i}h_{new}^{i},&\text{if$\phi\in(0^{\circ},\phi_{1})$}\\ 0,&\text{if $\phi\in[\phi_{1},\phi_{2}]$}\\ -\alpha^{i}h_{old}^{i}+\beta^{i}h_{new}^{i},&\text{if $\phi\in(\phi_{2},180^{\circ})$}\\ \end{cases}$ (11)
@@
We normalize the importance scores $\mathcal{I}_{\text{old}}=\{\mathcal{I}_{\text{old}}^{1},\dots,\mathcal{I}_{\text{old}}^{N}\}$ and $\mathcal{I}_{\text{new}}=\{\mathcal{I}_{\text{new}}^{1},\dots,\mathcal{I}_{\text{new}}^{N}\}$ independently to the range [0, 1]. This yields the final fusion weights $\alpha=\{\alpha^{1},\dots,\alpha^{N}\}$ and $\beta=\{\beta^{1},\dots,\beta^{N}\}$, which are then applied to the corresponding task vectors $h_{\text{old}}$ and $h_{\text{new}}$ for editing.
@@ 5 Experiments and Analysis
Dataset Method LLAMA2-7B LLAMA-7B
@@
LoRA 43.10 42.20 70.83 46.93 45.87 75.86
F-LearningLoRA 46.91 46.21 72.50 47.56 46.90 76.09
GeoEditLoRA 46.11 47.19 78.42 48.56 47.11 80.50
FT-c 49.02 46.96 67.37 47.33 45.51 68.14
Full-FT 81.02 74.67 70.51 70.52 66.69 65.26
MEND 29.77 25.86 71.54 30.99 27.12 69.83
@@
F-LearningFT 84.65 81.51 70.92 83.06 79.50 70.09
GeoEditFT 85.21 82.43 75.71 84.81 79.86 75.15
@@
LoRA 30.56 23.24 40.08 27.54 21.21 39.75
F-LearningLoRA 31.17 23.63 40.42 29.47 22.89 44.91
GeoEditLoRA 31.64 24.13 43.83 30.89 23.00 47.68
FT-c 29.23 19.32 19.70 26.97 17.90 20.09
Full-FT 65.99 44.08 28.34 32.13 31.95 32.51
MEND 14.77 14.67 90.93 17.51 16.27 89.64
@@
F-LearningFT 69.53 45.56 28.41 56.39 39.75 31.87
GeoEditFT 68.34 46.53 37.73 55.88 40.60 42.33
Table 1: Results on three metrics for the two datasets using LLAMA2-7B and LLAMA-7B.
@@
We use two widely recognized datasets: ZsRE Levy et al. (2017) and CounterFact (Meng et al., 2022). ZsRE is a Question Answering dataset that generates equivalence neighborhoods using question paraphrases produced through back-translation. CounterFact presents a more complex challenge with counterfactual data. We adopt the experimental setup from Yao et al. (2023), using the eval and edit subsets consisting of 19,085 and 10,000 examples, respectively. The datasets are partitioned into old and new knowledge categories, as in F-Learning Ni et al. (2024). For example, in ZsRE, old knowledge is modified to new knowledge, such as the change from “Los Angeles” to “New Orleans.” Further details and additional examples are provided in Appendix A.
@@
We evaluate GeoEdit against two types of methods: fine-tuning-based approaches, including full fine-tuning (Full-FT), LoRA Hu et al. (2021), FT-c Zhu et al. (2020), and F-Learning Ni et al. (2024); and locate-and-edit-based methods, including MEND Mitchell et al. (2022), ROME Meng et al. (2022), MEMIT Meng et al. (2023), and RECT Gu et al. (2024). Detailed descriptions are provided in the Appendix B.
@@
Following the setup of F-Learning Ni et al. (2024), we use LLAMA2-7B and LLAMA-7B as the base models and focus on updating old knowledge to new knowledge. First, we fine-tune the base model on the old knowledge for three epochs using full fine-tuning, resulting in the original model, which serves as the baseline for our experiments. At this stage, the model has fully learned the old knowledge, making it essential to analyze the angular divergence between the task vectors of the old and new knowledge. In GeoEdit, the encoder and decoder consists of 2-layer MLPs with dimensions $[4096\rightarrow 2048,2048\rightarrow 512]$ and $[512\rightarrow 2048,2048\rightarrow 4096]$, respectively, where $d_{latent}$ is set to 512. We set $\lambda$ in Eq. (9) to 0.5, $\phi_{1}$ and $\phi_{2}$ in Eq. (11) to $85^{\circ}$ and $95^{\circ}$, respectively. Further details on the experimental setup can be found in Appendix C.

### v2

@@ Abstract
Regular updates are essential for maintaining up-to-date knowledge in large language models (LLMs). However, existing training-based model editing methods often struggle to effectively incorporate new knowledge while preserving unrelated general knowledge. To address this challenge, we propose a novel framework called Geometric Knowledge Editing (GeoEdit). GeoEdit utilizes the geometric relationships of parameter updates from fine-tuning to differentiate between neurons associated with new knowledge updates and those related to general knowledge perturbations. By employing a direction-aware knowledge identification method, we avoid updating neurons with directions approximately orthogonal to existing knowledge, thus preserving the model’s generalization ability. For the remaining neurons, we integrate both old and new knowledge for aligned directions and apply a “forget-then-learn” editing strategy for opposite directions. Additionally, we introduce an importance-guided task vector fusion technique that filters out redundant information and provides adaptive neuron-level weighting, further enhancing model editing performance. Extensive experiments on two publicly available datasets demonstrate the superiority of GeoEdit over existing state-of-the-art methods.
@@ 1 Introduction
Large language models (LLMs) have demonstrated the ability to store vast amounts of knowledge during pre-training and retrieve it during inference Yao et al. (2023); Wang et al. (2024d). However, much of the knowledge in the real world is constantly evolving. For instance, the answer to the question “Who is the President of the United States?” was “Joe Biden” in 2024, but it is now “Donald Trump”. As a result, some knowledge that was once correct in LLMs can become obsolete or inaccurate Li and Chu (2024); Huang et al. (2024).
To address this issue, model editing methods have been developed to update the target new knowledge while preserving unrelated general knowledge within the model Hong and Lipani (2024); Ma et al. (2024); Wang et al. (2024e). Specifically, current model editing methods typically follow the “locate-and-edit” paradigm Wang et al. (2023, 2024a); Li et al. (2024). The core idea is first to locate influential weights in LLMs and then edit them by introducing a perturbation. Although effective, these approaches incur significant computational overhead to identify the important neurons and parameters Meng et al. (2022). Some methods also require sampling additional data (e.g., from Wikipedia) to mitigate the impact on the general knowledge within LLMs during editing Meng et al. (2023); Fang et al. (2024), introducing extra costs and potential biases.
@@
Additionally, to mitigate angular bias in high-dimensional space, GeoEdit employs a combined dimensionality reduction approach to more effectively extract angular information, ensuring the reliability of the edits. To optimize vector fusion, we introduce an importance-guided task vector fusion technique, which applies fine-grained weights to the vectors and suppresses noise from redundant parameters, further enhancing the effectiveness of model editing. Extensive experiments demonstrate that GeoEdit achieves the best performance among all fine-tuning-based methods and show its significant potential for complementing locate-and-edit methods, further enhancing performance.
@@ 2 Related Work
Knowledge editing has gained significant attention due to the increasing need to update the knowledge in LLMs Shengyuan et al. (2023); Wang et al. (2024d); Bi et al. (2024c, b). Existing methods can be classified into two main approaches:
Locate and edit methods usually locate influential parameters and then edit them by introducing a perturbation Zhang et al. (2024); Jiang et al. (2024); Xu et al. (2024). Classic methods like ROME Meng et al. (2022) and MEMIT Meng et al. (2023) use causal reasoning to identify key neuron activations and adjust specific weights. Additionally, Yu et al. (2023) employs gradient-based attribution to identify important weights. More recent approaches, such as AlphaEdit Fang et al. (2024), improve the method by projecting perturbations onto the null space of preserved knowledge, demonstrating strong performance.
Fine-tuning is an intuitive and straightforward way to update the model’s knowledge Feng et al. (2023); Gangadhar and Stratos (2024); Zheng et al. (2024). Recently, a series of PEFT methods, such as Prefix-Tuning Li and Liang (2021) and LoRA Hu et al. (2022), have made knowledge editing based on fine-tuning more feasible. Zhang et al. (2023) enhance update efficiency and adaptability by performing incremental parameter updates of varying magnitudes, which are determined by calculating the importance of the weight matrix.
@@
However, both of these methods struggle to balance new knowledge updates with preserving unrelated knowledge Gupta et al. (2024); Feng et al. (2024a); Chen et al. (2024). For instance, locate-and-edit methods typically require large additional datasets to capture general knowledge and avoid disruption during editing Wang et al. (2024c); Hsueh et al. (2024); Bi et al. (2024a); Zhang et al. (2025). Furthermore, locate-and-edit methods primarily focus on editing the MLP layers of the model, while fine-tuning methods offer the flexibility to adjust different regions of the model.
@@
Thus, our paper focuses on improving fine-tuning methods. By distinguishing between general knowledge and updated knowledge based on the angular divergence between the updated directions of old and new knowledge, our GeoEdit avoids updating general knowledge, ensuring model generalization while applying tailored strategies to enhance the effectiveness of knowledge updates. Additionally, our method can fine-tune parameters in regions different from those targeted by locate-and-edit methods, allowing for potential complementarity that further enhances performance.
3 Problem Statement
@@
Figure 2: Overview of GeoEdit. Step (a): Neuron-level task vectors $\tau_{old}$ and $\tau_{new}$ are extracted for both the old and new knowledge datasets using parametric arithmetic. Step (b): An auto-encoder is trained to project a low-dimensional representation of the task vectors, eliminating the angular bias issue in high-dimensional space. Step (c): The latent task vectors, $h_{new}$ and $h_{old}$, are reduced to two dimensions using t-SNE to compute the angular relationships, which are used to classify neurons based on the angle. Finally, after applying different editing strategies, we obtain the edited vector $\tau_{edit}$, which is added to the initial model to generate the edited model $f_{\theta_{e}}$.
@@ 4.1 Extracting Neuron-level Task Vectors
While prior research typically captures task vectors at the model level Ilharco et al. (2022), we propose extracting them at the neuron level for finer control. Let $\theta=\{\theta^{1},\theta^{2},\dots,\theta^{N}\}$ represent the $N$ neurons in the LLM, where the $i$-th neuron is represented by $\theta^{i}\in\mathbb{R}^{d_{n}}$ with $d_{n}$ dimensional parameters. The neuron-level task vectors are given by $\tau_{new}=\{\tau_{new}^{1},\tau_{new}^{2},\dots,\tau_{new}^{N}\}$, where $\tau_{new}^{i}$ corresponds to the new knowledge task vector for the $i$-th neuron  [^1] We define a “neuron” as the linear transformation corresponding to a single column in matrix $W\in\mathbb{R}^{in\times out}$, where $W$ consists of $out$ neurons.. This approach enables more granular analysis of parameter changes and selective editing of knowledge-specific neurons, enhancing model editing precision.
@@
4.2 Angular Relationship Extraction through Dimensionality Reduction
Due to the tendency of high-dimensional vectors to become nearly orthogonal, it is necessary to reduce the dimensionality of the original vectors in order to better capture the underlying angular relationships. However, experiments have shown that directly applying PCA or t-SNE for dimensionality reduction on $\tau$ yields suboptimal results. Therefore, we propose an alternative approach where an auto-encoder (AE) is first used to encode the high-dimensional vectors. This effectively filters out irrelevant information and extracts meaningful features. Subsequently, applying t-SNE to the encoded vectors allows for a more accurate representation of the true angular relationships.
Thus, we train a semantic encoder and decoder, both implemented using multi-layer perceptrons (MLPs). Specifically, the semantic encoder, denoted as SemEnc($\cdot$), maps the high-dimensional task vectors $\tau_{old}$ and $\tau_{new}$ into the latent space as:
@@ 4.3 Geometric Knowledge Editing
After training the AE, we project $\tau_{old}$ and $\tau_{new}$ into the latent space and then apply t-SNE to further project them into a 2D space to compute the angular relationships. GeoEdit then edit the original task vectors to obtain the edited task vector $\tau_{edit}$. This vector is subsequently added to the initial model, resulting in the final edited model $f_{\theta_{e}}$.
@@
For neuron $i$, we first use the encoder to reduce the dimensionality of $\tau_{old}^{i}$ and $\tau_{new}^{i}$ yielding the latent task vectors $h_{old}^{i}$ and $h_{new}^{i}$. We then apply t-SNE to obtain the 2D vectors $\hat{h}_{old}^{i}$ and $\hat{h}_{new}^{i}$. Next, we compute the angular divergence $\phi$ as:
$\phi=\arccos\frac{\hat{h}_{old}^{i}\cdot\hat{h}_{new}^{i}}{\left|\hat{h}_{old}^{i}\right|\cdot\left|\hat{h}_{new}^{i}\right|}$ (10)
@@
$\tau_{edit}^{i}=\begin{cases}\alpha^{i}\tau_{old}^{i}+\beta^{i}\tau_{new}^{i},&\text{if $\phi\in(0^{\circ},\phi_{1})$}\\ 0,&\text{if $\phi\in[\phi_{1},\phi_{2}]$}\\ -\alpha^{i}\tau_{old}^{i}+\beta^{i}\tau_{new}^{i},&\text{if $\phi\in(\phi_{2},180^{\circ})$}\\ \end{cases}$ (11)
@@
We normalize the importance scores $\mathcal{I}_{\text{old}}=\{\mathcal{I}_{\text{old}}^{1},\dots,\mathcal{I}_{\text{old}}^{N}\}$ and $\mathcal{I}_{\text{new}}=\{\mathcal{I}_{\text{new}}^{1},\dots,\mathcal{I}_{\text{new}}^{N}\}$ independently to the range [0, 1]. This yields the final fusion weights $\alpha=\{\alpha^{1},\dots,\alpha^{N}\}$ and $\beta=\{\beta^{1},\dots,\beta^{N}\}$, which are then applied to the corresponding task vectors $\tau_{\text{old}}$ and $\tau_{\text{new}}$ for editing.
@@ 5 Experiments and Analysis
Dataset Paradigm Method LLAMA2-7B LLAMA-7B
@@
Locate & edit MEND 29.77 25.86 71.54 30.99 27.12 69.83
@@
AlphaEdit 87.91 81.52 77.14 87.09 80.41 76.53
Fine-tuning LoRA 43.10 42.20 70.83 46.93 45.87 75.86
FT-c 49.02 46.96 67.37 47.33 45.51 68.14
Full-FT 81.02 74.67 70.51 70.52 66.69 65.26
F-Learning 84.65 81.51 70.92 83.06 79.50 70.09
GeoEdit 85.21 82.43 75.71 84.81 79.86 75.15
GeoEdit 88.13 82.07 79.75 87.76 80.70 77.98
@@
Locate & edit MEND 14.77 14.67 90.93 17.51 16.27 89.64
@@
AlphaEdit 71.79 48.36 36.07 60.01 38.19 41.70
Fine-tuning LoRA 30.56 23.24 40.08 27.54 21.21 39.75
FT-c 29.23 19.32 19.70 26.97 17.90 20.09
Full-FT 65.99 44.08 28.34 32.13 31.95 32.51
F-Learning 69.53 45.56 28.41 56.39 39.75 31.87
GeoEdit 68.34 46.53 37.73 55.88 40.60 42.33
GeoEdit 72.20 48.57 38.71 60.89 41.37 43.99
Table 1: Results on three metrics for the two datasets using LLAMA2-7B and LLAMA-7B. The best-performing method for each paradigm is highlighted in bold. AlphaEdit and our GeoEdit each achieve the best performance within their respective paradigms. Notably, the optimal performance is attained by GeoEdit, which results from applying GeoEdit to the non-located parameters in AlphaEdit, effectively combining the strengths of both methods.
@@
We use two widely recognized datasets: ZsRE Levy et al. (2017) and CounterFact (Meng et al., 2022). We adopt the experimental setup from Yao et al. (2023), using the eval and edit subsets consisting of 19,085 and 10,000 examples, respectively. The datasets are partitioned into old and new knowledge categories, as in F-Learning Ni et al. (2024). For example, in ZsRE, old knowledge is modified to new knowledge, such as the change from “Los Angeles” to “New Orleans.” Further details and additional examples are provided in Appendix A.
@@
We evaluate GeoEdit against two types of methods: fine-tuning-based approaches, including full fine-tuning (Full-FT), LoRA Hu et al. (2021), FT-c Zhu et al. (2020), and F-Learning Ni et al. (2024); and locate-and-edit-based methods, including MEND Mitchell et al. (2022), ROME Meng et al. (2022), MEMIT Meng et al. (2023), RECT Gu et al. (2024) and AlphaEdit Fang et al. (2024). Detailed descriptions are provided in the Appendix B.
@@
Following the setup of F-Learning, we first fine-tune the base model on the old knowledge for three epochs, resulting in the original model, which serves as the baseline for our experiments. In GeoEdit and F-Learning, we use LoRA to enhance the efficiency of fine-tuning. The encoder and decoder consists of 2-layer MLPs with dimensions $[4096\rightarrow 2048,2048\rightarrow 512]$ and $[512\rightarrow 2048,2048\rightarrow 4096]$, respectively, where $d_{latent}$ is set to 512. We set $\lambda$ in Eq. (9) to 0.5, $\phi_{1}$ and $\phi_{2}$ in Eq. (11) to $85^{\circ}$ and $95^{\circ}$, respectively. Further details on the experimental setup can be found in Appendix C.